OpenStudy (anonymous):

Consider the line that passes through the points (-5, 9) and (-2, 0). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 9 = -1/3(x + 5) y - 9 = -3(x + 5) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (cwrw238):

slope of the line = 0-9 / -2-(-5) = -9/3 and the line passes through (-5,9) plug theses values into the general equation y - y1 = m(x - x1) where m = slope and x1 = -5 and y1 = 9

OpenStudy (anonymous):

so a ?

OpenStudy (cwrw238):

the slope = m = -9/3 = -3 right? so how can it be a?

OpenStudy (anonymous):

Only equation #2 represents the line that passes through the two given points

OpenStudy (cwrw238):

yes